# A Survey on Inference Optimization Techniques for Mixture of Experts Models

### v1

@@ 4.1. Expert Parallelism
Expert parallelism is a key technique for deploying large MoE models across multiple devices to enable distributed execution. As Figure 7-(a) shown (10.1145/3627703.3650083, ), when distributing MoE layers, each device holds a subset of experts along with all other non-expert parameters. In special cases where an expert is very large, each device may contain only one expert. During the execution of an MoE layer, inputs are processed on each device by the attention module and gate module. Subsequently, an all-to-all communication operation redistributes tokens to corresponding devices based on the gate output, allowing each expert to process its assigned inputs. Finally, another all-to-all communication operation occurs to exchange the experts’ outputs and send them back to the originating devices. Consequently, the execution time of MoE layers is dominated by two primary phases: computation and communication. To accelerate inference, various studies have proposed optimizations for these phases.
Existing research focuses on optimizing MoE training and inference performance along four main dimensions: parallel strategies, load balancing, all-to-all communication, and task scheduling.
@@ 4.1.1. Parallelism Strategy Design
In addition to relying solely on expert parallelism (he2021fastmoe, ; lepikhin2020gshard, ), recent studies combine multiple parallel strategies to maximize parallelism and achieve more efficient distributed computing. Tutel (hwang2023tutel, ) dynamically switches parallel strategies at each iteration without incurring extra overhead. MoESys (10528887, ) employs an elastic MoE training strategy with 2D prefetching and fusion communication over hierarchical storage for efficient parallelism. Alpa (Alpa, ) automatically derives efficient parallel execution plans at both inter-operator and intra-operator levels. DeepSpeed-TED (DeepSpeedted, ) implements a hybrid parallel algorithm combining data, tensor, and expert parallelism. BaGuaLu (BaGuaLu, ) integrates MoE and data parallelism to scale models effectively. SmartMoE (288691, ) introduces a two-stage approach to identify optimal parallel strategies for data-sensitive models.
@@ 4.1.2. Load Balancing
Token processing in MoE models can lead to imbalances, where some experts receive significantly more tokens than others. This disparity results in some devices experiencing high overheads while others remain idle, ultimately delaying overall computation. To address this issue, some works propose to add auxiliary loading balancing loss to train the gate module (shazeer2017outrageously, ; fedus2022switch, ; lepikhin2020gshard, ), while other works implement gate modules with special hash functions to avoid imbalance loading (roller2021hash, ). Advanced techniques also focus on expert placement optimization. For example, Prophet (10319949, ) leverages temporal locality to fine-tune resource allocation and expert placement, while MoE-Prediction (cong2024prediction, ) monitors expert loads and adapts placements based on transient and stable states. Lazarus (wu2024lazarus, ) allocates expert replicas using an optimal placement algorithm, and FlexMoE (nie2023flexmoe, ) employs fine-grained replication strategies to dynamically adjust the expert-to-device mapping at runtime. Additional strategies, such as those proposed by MoE-Deploy (huang2023towards, ) and Brainstorm (Brainstorm, ), rely on historical allocation data and expert reordering to balance loads. Furthermore, techniques like Lynx (gupta2024lynx, ) reduce the number of active experts during batch inference, and BaseLayers (lewis2021base, ) formulates token-to-expert allocation as a linear assignment problem to ensure equitable token distribution. And MoE-ECR (moeecr, ) allows experts to select the top-k tokens instead of letting each token choose the top-k experts, ensuring that each expert has a fixed bucket size.
@@ 4.1.3. All-to-all Communication Optimization
All-to-all communication is a significant bottleneck in MoE layer execution (li2024optimizing, ; 10.1145/3503221.3508418, ; DeepSpeedmoe, ), prompting substantial research to optimize this operation. Many efforts focus on hierarchical communication strategies and data compression techniques to reduce overhead. For instance, Tutel (hwang2023tutel, ) introduces a two-dimensional hierarchical all-to-all algorithm, while HetuMoE (nie2022hetumoe, ) and DeepSpeed-MoE (DeepSpeedmoe, ) optimize communication by combining intra-node and inter-node hierarchies. Data compression strategies, such as those employed by DeepSpeed-TED (DeepSpeedted, ) and TA-MoE (chen2022ta, ), minimize data transfer by eliminating unnecessary information and adapting data volume to the network topology. Innovative paradigms like Janus (10.1145/3603269.3604869, ) adopt a data-centric approach, where experts are moved between devices instead of tokens, significantly reducing communication costs. Similarly, ExFlow (yao2024exploiting, ) exploits inter-layer expert affinity to halve the number of all-to-all operations per MoE layer, and Aurora (li2024optimizing, ) optimizes token transmission to avoid bandwidth contention. LocMoE (li2024locmoe, ) and Parm (10621327, ) further enhance communication efficiency by converting inter-node communication to intra-node operations and overlapping intra-node with inter-node communications. And Lina (288705, ) accelerates the all-to-all communication by prioritizing al-to-all operation over all-reduce operation.
@@ 4.1.4. Task Scheduling
To overlap communication and computation tasks to reduce end-to-end runtime, various task scheduling strategies have been developed. Advanced scheduling strategies leverage architectural and algorithmic innovations to achieve this overlap effectively. For example, ScMoE (cai2024shortcut, ) integrates shortcut-connected MoE architecture to enable simultaneous communication and computation, while HiDup (HiDup, ) uses microbatching to alternate between computation and communication within a single training step. ScheMoE (10.1145/3627703.3650083, ) provides a generic framework for optimizing task scheduling, and PipeMoE (10228874, ) pipelines tasks to hide communication latency through a performance model. Additionally, EPS-MoE (qian2024eps, ) dynamically selects optimal kernel implementations to adaptively overlap FFN module computation with all-to-all communication, further improving runtime efficiency.
@@ 4.2.1. Expert Prefetching
To minimize waiting time for the required experts, some studies focus on optimizing prefetching techniques, which overlap expert loading with GPU computation. Several works predict the experts needed for subsequent layers. For example, MoE-Offloading (eliseev2023fast, ), AdapMoE (zhong2024adapmoe, ), and HOBBIT (tang2024hobbit, ) use current gating inputs as predictors for the next-layer gating inputs, prefetching the required experts in advance. Pre-gated MoE (hwang2024pre, ) introduces a pre-gated MoE structure that identifies the experts required by the next layer during the computation of the current layer and preloads them accordingly. EdgeMoE (yi2023edgemoe, ) constructs a prediction table using a calibrated dataset, leveraging the current layer experts to predict the next layer’s experts. DyNN-Offload (dynnoffload, ) trains a pilot model for each layer to predict the required experts for the next layer. MoE-Infinity (xue2024moe, ) tracks the request-level usage frequency of each expert and prefetches high-priority experts based on this frequency. ProMoE (song2024promoe, ) uses a learned predictor to prefetch experts in a sliding-window manner. Additionally, some works aim to predict all the required experts in the current forward pass in a single step. For instance, ExpertFlow (he2024expertflow, ) trains a predictor to anticipate all the required experts for the current forward pass and prefetches them in advance. SiDA (du2024sida, ) employs a hash function to predict activated experts for each token in the sequence based on its embedding. Similarly, Read-ME (readmoe, ) constructs a pre-gating router decoupled from the MoE backbone to predict all required experts at once.
@@ 4.2.2. Expert Caching
Since GPU memory can hold only a subset of experts (expert cache), optimizing the cache management policy is critical to improving the cache hit ratio and reducing transfer time. Several works (eliseev2023fast, ; yi2023edgemoe, ; hwang2024pre, ; yuan2024efficient, ; huang2023towards, ) adopt the Least Recently Used (LRU) policy, as recently used experts are more likely to be accessed again. In contrast, MoE-Infinity (xue2024moe, ) employs a Least Frequently Used (LFU) policy due to its request-level tracking capability. Fiddler (kamahori2024fiddler, ) maintains the most critical experts in GPU memory using a static dataset-based approach, whereas SwapMoE (kong2023serving, ) dynamically updates the most important experts. AdapMoE (zhong2024adapmoe, ) introduces a dynamic cache size for different model layers, driven by its adaptive gating algorithm. Meanwhile, HOBBIT (tang2024hobbit, ) proposes a multidimensional cache policy combining LRU, LFU, and a novel Least High Precision Used (LHU) strategy for its mixed-precision expert cache.

### v2

@@ 4.1. Expert Parallelism
Expert parallelism is a key technique for deploying large MoE models across multiple devices to enable distributed execution. As Figure 7-(a) shown, when distributing MoE layers, each device holds a subset of experts along with all other non-expert parameters (10.1145/3627703.3650083, ). In special cases where an expert is very large, each device may contain only one expert. During the execution of an MoE layer, inputs are processed on each device by the attention module and gate module. Subsequently, an all-to-all communication operation redistributes tokens to corresponding devices based on the gate output, allowing each expert to process its assigned inputs. Finally, another all-to-all communication operation occurs to exchange the experts’ outputs and send them back to the originating devices. Consequently, the execution time of MoE layers is dominated by two primary phases: computation and communication. To accelerate inference, various studies have proposed optimizations for these phases.
Existing research focuses on optimizing MoE training and inference performance along four main dimensions: parallelism strategy, load balancing, all-to-all communication, and task scheduling.
@@ 4.1.1. Parallelism Strategy Design
In addition to relying solely on expert parallelism (he2021fastmoe, ; lepikhin2020gshard, ), recent studies combine multiple parallel strategies to maximize parallelism and achieve more efficient distributed computing. Tutel (hwang2023tutel, ) dynamically switches parallelism strategies at each iteration without incurring extra overhead. Specifically, this work employs a single distribution layout that encompasses all optimal strategies, thereby eliminating the need to reformat input data or weights when switching parallelism strategies. Alpa (Alpa, ) reclassifies conventional data and model parallelism into intra-operator and inter-operator parallelism based on the key observation that different parallelization techniques have varying bandwidth requirements for communication. In typical clusters, co-located devices have high communication bandwidth, whereas distant devices have limited communication bandwidth. With this reclassification, Alpa can automatically derive efficient parallel execution plans that incorporate data, operator, and pipeline parallelism based on the model description and cluster configuration. DeepSpeed-TED (DeepSpeedted, ) implements a hybrid parallel algorithm combining data, tensor, and expert parallelism. It designs a tiled version of the optimizer to reduce peak memory consumption arising from the combination of these three parallelism forms, thereby enabling the training of MoE models with 4–8× larger base models than the current state-of-the-art. BaGuaLu (BaGuaLu, ) proposes a parallel strategy called MoDa that combines MoE parallelism and data parallelism to effectively scale models. Specifically, MoDa places data parallelism within a single supernode while maintaining MoE communication globally across supernodes. This approach is based on the observation that All-reduce communication (in data parallelism) incurs much higher costs than All-to-all communication (in MoE parallelism) during the training of its target large-scale models. SmartMoE (288691, ) explores the space of hybrid parallelism with an awareness of heterogeneous workloads, incorporating potentially faster candidate parallel strategies. It introduces a two-stage approach to identify optimal parallel strategies for data-sensitive models, periodically determining whether a new parallel strategy should be employed at runtime. MPMoE (10494556, ) proposes a lightweight profile-based algorithm that leverages profiling information to identify the most suitable configuration at runtime, optimizing pipeline parallelism. Additionally, in serverless computing, MoE-SLC (liu2025optimizing, ) proposes a Bayesian optimization framework with multi-dimensional $\delta$-greedy search to learn expert selections for expert parallelism and optimize the MoE deployment to achieve optimal billed cost.
@@ 4.1.2. Load Balancing
Token processing in MoE models can lead to imbalances, where some experts receive significantly more tokens than others. This disparity results in certain devices experiencing high overheads while others remain idle, ultimately delaying overall computation. To address this issue, some works propose adding auxiliary load balancing loss terms to train the gate module (shazeer2017outrageously, ; fedus2022switch, ; lepikhin2020gshard, ), while others implement gate modules with special hash functions to prevent load imbalance (roller2021hash, ). Advanced techniques focus on optimizing expert placement. For example, Prophet(10319949, ) builds a load balance performance model to estimate the execution time of the MoE layer for each expert placement and then uses a greedy search algorithm to find a well-balanced expert placement. MoE-Prediction(cong2024prediction, ) traces and analyzes the loads of each expert during the training process, then employs classical prediction algorithms to forecast the expert load proportions, providing valuable guidance for expert placement in MoE training. Lazarus(wu2024lazarus, ) allocates expert replicas using an optimal placement algorithm that assigns more replicas to popular experts to help balance the load. FlexMoE(nie2023flexmoe, ) employs fine-grained replication strategies that select specific heavy experts and duplicate them across multiple devices, dynamically adjusting the expert-to-device mapping at runtime to address routing imbalances. Additional strategies, such as those proposed by MoE-Deploy(huang2023towards, ) and Brainstorm(Brainstorm, ), rely on historical allocation data and expert reordering to balance loads. Furthermore, techniques like Lynx(gupta2024lynx, ) reduce the number of active experts during batch inference, and BaseLayers(lewis2021base, ) formulates token-to-expert allocation as a linear assignment problem to ensure equitable token distribution. Additionally, MoE-ECR (moeecr, ) allows experts to select the top-k tokens instead of letting each token choose the top-k experts, ensuring that each expert has a fixed bucket size.
@@ 4.1.3. All-to-all Communication Optimization
All-to-all communication is a significant bottleneck in MoE layer execution (li2024optimizing, ; 10.1145/3503221.3508418, ; DeepSpeedmoe, ), prompting substantial research aimed at optimizing this operation. Many efforts focus on hierarchical communication strategies and data compression techniques to reduce overhead. For instance, Tutel(hwang2023tutel, ), HetuMoE(nie2022hetumoe, ), and DeepSpeed-MoE (DeepSpeedmoe, ) utilize the hierarchical all-to-all algorithm to optimize communication by combining intra-node and inter-node hierarchies. Compared to traditional all-to-all operations, which allow all GPUs to communicate directly with one another, this hierarchical approach first gathers data from all GPUs within a node to a single GPU (intra-node) and then communicates between nodes (inter-node). This method speeds up communication by fully utilizing both levels of bandwidth. Data compression strategies, such as those employed by DeepSpeed-TED(DeepSpeedted, ) and TA-MoE(chen2022ta, ), minimize data transfer by eliminating unnecessary information and adapting data volume to the network topology. Innovative paradigms like Janus (10.1145/3603269.3604869, ) adopt a data-centric approach, moving experts between devices instead of tokens, thereby significantly reducing communication costs. ExFlow(yao2024exploiting, ) reduces the number of all-to-all operations from two to one by ensuring token context coherence, meaning that all GPUs have the context of all requests rather than only their own. It then exploits inter-layer expert affinity to further reduce the number of all-to-all operations by optimally placing experts in each layer, increasing the probability that tokens remain on local GPUs. Aurora(li2024optimizing, ) orders token transmission to avoid bandwidth contention, achieving minimal all-to-all communication costs through theoretical derivation. LocMoE(li2024locmoe, ) and Parm(10621327, ) further enhance communication efficiency by converting inter-node communication to intra-node operations and overlapping intra-node with inter-node communications. Additionally, Lina (288705, ) accelerates all-to-all communication by prioritizing all-to-all operations over all-reduce operations.
@@ 4.1.4. Task Scheduling
To overlap communication and computation tasks and reduce end-to-end runtime, various task scheduling strategies have been developed. Advanced scheduling strategies leverage architectural and algorithmic innovations to achieve this overlap effectively. For example, ScMoE (cai2024shortcut, ) introduces a shortcut-connected MoE architecture that processes representations from both the current and preceding layers, rather than solely processing representations from the current layer. Specifically, ScMoE employs a top-1 MoE module to handle representations from the preceding layer via a shortcut connection, while a shared expert processes the current layer’s representations. This structure effectively decouples communication from its conventional sequence, thereby enhancing the overlap between communication and computation. HiDup (HiDup, ) splits the input data on each device into two microbatches with no dependencies between them. Consequently, the training of the two microbatches on each device can be carried out in parallel, effectively overlapping the computation of one microbatch with the communication of the other. MoESys (10528887, ) employs an elastic MoE training strategy with 2D prefetching and fusion communication over hierarchical storage to overlap computation with the parameter readiness from the hierarchical storage. ScheMoE(10.1145/3627703.3650083, ) first modularizes the computing tasks (data compression and expert computation) and communication tasks (collective communication). It then designs an adaptive optimal scheduling algorithm to pipeline the communication and computing tasks based on these modularized operations. Similarly, PipeMoE(10228874, ) designs a performance model to predict the computation and communication costs for various MoE workloads and then proposes an optimal polynomial-time solution to pipeline tasks, thereby hiding communication latency based on the performance model’s results. Additionally, EPS-MoE (qian2024eps, ) dynamically selects optimal kernel implementations to adaptively overlap FFN module computation with all-to-all communication, further improving runtime efficiency.
DeepSpeed-MoE (DeepSpeedmoe, ) FasterMoE (10.1145/3503221.3508418, )
Method Speedup Metric Method Speedup Metric
HetuMoE (nie2022hetumoe, ) 5.88x Time per iteration TA-MoE (chen2022ta, ) 1.39x Tokens per second
TA-MoE (chen2022ta, ) 1.31x Tokens per second PipeMoE (10228874, ) 1.69x Time per iteration
FlexMoE (nie2023flexmoe, ) 1.70x Time to converge FlexMoE (nie2023flexmoe, ) 1.30x Time to converge
Prophet (10319949, ) 2.39x Time per iteration Prophet (10319949, ) 1.43x Time per iteration
Parm (10621327, ) 3.00x Time per iteration ScheMoE (10.1145/3627703.3650083, ) 1.25x Time per iteration
Lazarus (wu2024lazarus, ) 3.40x Tokens per second MPMoE (10494556, ) 1.60x Time per iteration
DeepSpeed-TED (DeepSpeedted, ) 1.35x Time per iteration SMARTMOE (288691, ) 1.88x Time per iteration
Table 5. Speedup of parallelism systems when compared to DeepSpeed-MoE and FasterMoE in MoE training.
Method Speedup Metric Baseline Optimized Techniques
Lina (288705, ) 1.63x Time per batch DeepSpeed (deepspeed, ) Prioritize all2all over allreduce
TUTEL (hwang2023tutel, ) 2.03x Time per batch Fairseq (fairseq, ) Design an identical layout for weights
Aurora (li2024optimizing, ) 1.81x Time per batch Lina (288705, ) Strategically order token transmissions
ExFlow (yao2024exploiting, ) 2.2x Tokens per second Deepspeed-MoE (DeepSpeedmoe, ) Reduce two all2all operations into one
Brainstorm (Brainstorm, ) 3.33x Time per batch Tutel (hwang2023tutel, ) Preload weight with profiling statics
MoE-Deploy (huang2023towards, ) 3.32x Tokens per second Tutel (hwang2023tutel, ) Combine high-used experts with lows
DeepSpeed-MoE (DeepSpeedmoe, ) 7.25x Time per batch PyTorch (pytorch, ) Groupe tokens with same data path
Table 6. Speedup of parallelism systems in MoE Inference.
Method Benefits Baseline
Lina (288705, ) Gain 2.3x reduction in all2all communication DeepSpeed (deepspeed, )
Prophet (10319949, ) Improve the load balance by 1.75x to 12.06x FasterMoE (10.1145/3503221.3508418, )
ScheMoE (10.1145/3627703.3650083, ) Gain 1.4x to 2x speedup in all2all communication NCCLA2A (nie2022hetumoe, ), 2DH-A2A (hwang2023tutel, )
MPMoE (10494556, ) Achieve a 53% decrease in memory usage FasterMoE (10.1145/3503221.3508418, )
MoESys (10528887, ) Achieve a 18% decrease in memory usage DeepSpeed (deepspeed, )
Aurora (li2024optimizing, ) Achieve a 1.57x to 1.72x increase in GPU utilization Lina (288705, )
MoE-Deploy (huang2023towards, ) Achieve a 79.6% decrease in memory usage Tutel (hwang2023tutel, )
DeepSpeed-TED (DeepSpeedted, ) Support 1.09x to 4.8x larger MoE models DeepSpeed-MoE (DeepSpeedmoe, )
Table 7. Additional benefits of existing systems.
4.1.5. Performance Summarization
To assess the performance of the methods discussed above, we extract relevant data from their respective papers. Table 5 shows the training speedup of various systems relative to DeepSpeed-MoE and FasterMoE, both of which are widely recognized by researchers and serve as excellent starting points for related studies. Although different systems evaluate their methods using various configurations, a rough comparison can still be made due to the shared baselines. For instance, HetuMoE demonstrates exceptional performance, with the highest speedup compared to other systems. In addition to training speedup, we also report inference speedup and highlight the main optimization techniques of several systems in Table 6. However, comparing performance across systems is challenging due to differences in their baselines. Beyond speedup, some systems offer additional benefits, such as reduced memory usage. As shown in Table 7, many systems focus on optimizing memory usage, which is critical given the large size of MoE models. Additionally, some studies also optimize GPU utilization and communication efficiency.
@@ 4.2.1. Expert Prefetching
To minimize the waiting time for the required experts, some studies focus on optimizing prefetching techniques that overlap expert loading with GPU computation. Many works predict the experts needed for subsequent layers. For example, Mixtral-Offloading (eliseev2023fast, ), AdapMoE (zhong2024adapmoe, ), and HOBBIT (tang2024hobbit, ) use current gating inputs to feed the gating module for the following layers, predicting the required experts and prefetching them in advance. This prediction is based on the key observation that the inputs to the gating function in the MoE module share high similarity across successive layers, due to the residual structure of LLMs. As reported in these works, the prediction accuracy can reach about 90%, significantly enhancing the effectiveness of this approach. Pre-gated MoE (hwang2024pre, ) introduces a pre-gated MoE structure, which identifies the experts required by the next layer during the computation of the current layer and preloads them accordingly. Although this new structure is suitable for prefetching, it may cause a decrease in model accuracy. EdgeMoE (yi2023edgemoe, ) constructs a prediction table using a calibrated dataset, leveraging the experts from the current layer to predict the experts for the next layer. This table is built on the observation that the expert activations of sequential layers are statistically correlated. This means that given prior knowledge of the expert activations from layers 1 to n-1, we can estimate the activation probability of each expert at the n-th layer with high confidence. DyNN-Offload (dynnoffload, ) trains a pilot model for each layer to predict the required experts for the next layer. MoE-Infinity (xue2024moe, ) tracks the request-level usage frequency of each expert and prefetches high-priority experts based on this frequency. It prefetches experts across multiple layers concurrently, rather than just the next layer, but prioritizes the experts closest to the currently executed layer. ProMoE (song2024promoe, ) uses a learned predictor to prefetch experts in a sliding-window manner. This predictor, a small neural network (a two-layer MLP), learns the correlation between layer inputs and expert selections. It can use the inputs of the i-th layer to predict the expert selections of the (i+k)-th layer, where k is the sliding window size (e.g., 4 or 8). ProMoE runs LLM inference for multiple iterations to collect training data for the predictor in the offline stage and runs the predictor on the CPU to hide its execution latency by overlapping it with the LLM inference during the online stage. Additionally, some works aim to predict all the required experts for the current forward pass at the start of the forward iteration. For example, ExpertFlow (he2024expertflow, ) trains a transformer-based predictor, which is particularly well-suited for this task due to its exceptional capabilities in contextual modeling, to predict all the required experts for the current forward pass and prefetch them in advance. This predictor is trained as a classification task, outputting the state of each expert as either active or idle. SiDA (du2024sida, ) employs a network-based hash function to predict the activated experts for each token at each layer for the current incoming batch of data. This predictor is constructed with an LSTM and trained using truncated knowledge distillation. Similarly, Read-ME (readmoe, ) constructs a pre-gating router decoupled from the MoE backbone to predict all the required experts at once.
@@ 4.2.2. Expert Caching
Since GPU memory can only hold a subset of experts (the expert cache), optimizing the cache management policy is crucial for improving the cache hit ratio and reducing transfer time. Several works (eliseev2023fast, ; yi2023edgemoe, ; hwang2024pre, ; yuan2024efficient, ; huang2023towards, ) adopt the Least Recently Used (LRU) policy, based on the assumption that recently used experts are more likely to be accessed again. According to Mixtral-8x7B (jiang2024mixtral, ), when an expert is selected for the i-th token, it has a significantly higher probability (more than 10% in most layers) of being selected again for the (i+1)-th token compared to a random selection. In contrast, MoE-Infinity (xue2024moe, ) employs a variant of the Least Frequently Used (LFU) policy to define the caching priority of experts, leveraging its request-level tracking capability. Fiddler (kamahori2024fiddler, ) maintains a set of important experts in GPU memory using a static dataset to profile the active times of experts, treating these times as a measure of their importance. This method demonstrates an improvement of approximately 3 percentage points in hit rate compared to random placement. Similarly, SwapMoE (kong2023serving, ) stores a set of important experts in GPU memory based on a defined expert importance score, but it dynamically updates the cached experts at a suitable frequency. AdapMoE (zhong2024adapmoe, ) introduces a dynamic cache size for different model layers, which is driven by its adaptive gating algorithm. This algorithm selects a varying number of experts for different layers, thus adjusting the cache size accordingly. Specifically, it assigns a larger cache size to layers that select more experts or experience lower prediction accuracy. HOBBIT (tang2024hobbit, ) proposes a multidimensional cache policy that combines LRU, LFU, and a novel Least High Precision Used (LHU) strategy for its mixed-precision expert cache. Since high-precision experts incur higher loading latencies than low-precision experts, the LHU policy prioritizes high-precision experts to reduce loading costs. Additionally, CacheMoE (skliar2024mixture, ) employs a cache-aware strategy that adjusts the router logits $z=G(x)$ to enhance the likelihood of selecting experts already present in the cache. This approach effectively increases the expert cache hit ratio, thereby reducing the costs associated with expert loading.